OpenStudy (anonymous):

To travel 60 miles, it takes sue, riding a Harley, 2 hours less time than it takes Matt to travel 50 miles riding a bicycle. Sue travels 10 miles per hour faster than Matt. Find the times and rates of Matt and sue.

OpenStudy (wolfe8):

Welcome to OpenStudy. I think this is more math or physics but rate will be distance over time. Let SR be Sue's rate, ST be Sue's time, MR be Matt's rate, and MT be Matt's time. So you have:- SR= 60/ST ST = MT - 2 MR = 50/MT SR = MR + 10 I think I have them all.

OpenStudy (anonymous):

omg thanks so much! Can you explain how you did that?

OpenStudy (wolfe8):

I basically translated the sentences into math sentences. Do you know how to do that?

OpenStudy (anonymous):

not really.. but I kind of understand how you did that. this is an open ended problem that needed me to show work, is there work that I could put with that too or is it basically mental math?

OpenStudy (anonymous):

and also, you put SR (susie's rate) twice.. I'm confused why? and how come there is no rate put for matt?

OpenStudy (wolfe8):

I think mental math or maybe it's called logic. I'd say it's just a skill you can learn. Well one equation gives us ST and the other gives MR. All we know for MR is he traveled 50 miles in MT using the formula of speed(rate)=distance/time

OpenStudy (anonymous):

okay thanks so much for your help! are you like 100% positive that this is right? If not is there another categorie I could place this under to get more help?

OpenStudy (wolfe8):

You're very welcome. Yea you can put this in Math. Click on Find More Subjects at the top of this page. I believe I am right unless I made a careless mistake. I'd like to welcome you to OpenStudy. Do take time to read our CoC: http://openstudy.com/code-of-conduct and enjoy your stay. If you are satisfied with any question, please reward a medal to whoever you think deserves it and close the question. We also have a fan and testimony system. See you around :)

OpenStudy (anonymous):

Thanks! :) How do you give a medal!?

OpenStudy (wolfe8):

You just did haha. Click on Best Response at the right of a person's reply.

OpenStudy (wolfe8):

And thanks for the medal :)

OpenStudy (anonymous):

Thank youuu!